OpenStudy (anonymous):

Find the point that lies on the line described by the equation below. y - 8 = -3(x - 6) A. (18, 8) B. (6, 8) C. (8, 6) D. (6, -8)

OpenStudy (amistre64):

x=6; y=8

OpenStudy (anonymous):

B

OpenStudy (amistre64):

y - Py = m(x - Px) (Px,Py)

OpenStudy (llort):

don't just give the answer, help him out... y=-3x+18+8 y=-3x+26 from there we can get the points: (0,26) (1,23) (2,20) (3,17) (4,14) (5,11) ...